# Beyond the Black Box: Theory and Mechanism of Large Language Models

### v1

@@ 1 Introduction
The recent emergence of Large Language Models (LLMs) has marked a profound paradigm shift in the landscape of Artificial Intelligence (AI). Models such as ChatGPT [openai2022introducingchatgpt], DeepSeek [guo2025deepseek], Qwen [bai2023qwen], Llama [touvron2023llama], Gemini [team2023gemini], and Claude [caruccio2024claude] have transcended the boundaries of traditional Natural Language Processing (NLP) [transformer], demonstrating capabilities that impact nearly every facet of modern society. As these systems scale, they exhibit behaviors that mimic human-like reasoning [CoT], sparking a global transformation in how we interact with information.
In the history of technological development, engineering triumphs are often inextricably linked to scientific innovation. However, the synchronization between theory and application is rarely instantaneous. Consider the trajectory of nuclear physics: from Einstein’s formulation of the mass-energy equivalence equation ($E=mc^{2}$) in 1905 [einstein1905does] to the detonation of the first atomic bomb at Los Alamos in 1945 [rhodes2012making], scientists and engineers traversed a forty-year journey to translate theoretical insight into physical reality [hoddeson1993critical]. A similarly extended timeline defines the current era of AI. Approximately 33 years elapsed between the proposal of the Universal Approximation Theorem [hornik1989multilayer], which provided the mathematical assurance that neural networks could represent any continuous function, and the emergence of ChatGPT [openai2022introducingchatgpt], the definitive proof of that potential. Looking ahead from our current vantage point, the quest for Artificial General Intelligence (AGI) necessitates a balanced synergy where continuous theoretical research and rigorous engineering implementation are recognized as equally indispensable pillars.
@@
Throughout these decades, researchers have relentlessly pursued the essence of intelligence through diverse engineering and scientific lenses. At this pivotal moment, with the empirical success of LLMs, we appear closer than ever to unveiling the nature of intelligence. Yet, a paradox persists within our current standing: despite the monumental engineering successes of LLMs, our theoretical understanding of them remains disproportionately nascent. While deep learning theory has advanced substantially [roberts2022principles], the specific phenomena emerging from LLMs loom like a “dark cloud” over the field, shattering previous intuitions and challenging established statistical learning paradigms [scalinglaw1]. Consequently, we are currently forced to treat LLMs largely as “black boxes” [linardatos2020explainable, zhao2024explainability]. They function exceptionally well, yet their internal mechanisms of operation, the how and why behind their efficacy, remain elusive.
@@
The difficulty in piercing this black box stems primarily from two dimensions. First, the sheer scale of LLMs introduces unprecedented complexity [scalinglaw1, scalinglaw2]. With parameter counts reaching the trillions and a natural language state space that is combinatorially vast, accurately analyzing the learning dynamics and optimization landscape becomes an arduous mathematical challenge. Second, LLMs exhibit numerous “emergent” phenomena that do not appear in smaller models, such as hallucination [xu2024hallucination], in-context learning (ICL) [brown2020language], scaling laws [scalinglaw1], and sudden “aha moments” during training [guo2025deepseek]. These phenomena are difficult to unify under a single theoretical framework, rendering the modeling of LLMs a fragmented endeavor. Consequently, current analyses of LLM theory and mechanisms are often scattered, isolated within specific sub-topics without a holistic view.
@@
To address this fragmentation, this survey proposes a comprehensive, lifecycle-based perspective. Following the standard LLM pipeline, we categorize the theoretical landscape into six distinct stages as illustrated in fig. 1: the Data Preparation Stage, Model Preparation Stage, Training Stage, Alignment Stage, Inference Stage, and Evaluation Stage. By categorizing popular topics and theoretical advances into these stages, we aim to provide a structured roadmap that connects empirical observations with their underlying mechanisms.
The main contributions of this work are as follows:
  1. 1.
    A Unified Lifecycle-Based Taxonomy. We propose a structural framework that organizes the fragmented landscape of LLM research into six distinct stages: Data Preparation, Model Preparation, Training, Alignment, Inference, and Evaluation. This lifecycle perspective allows for a systematic exploration of the theoretical underpinnings at every step of an LLM’s development.
  2. 2.
    Systematic Review of Theory and Mechanisms. Moving beyond engineering heuristics, we provide a comprehensive review of the foundational theories and internal mechanisms driving LLMs. We analyze core theoretical issues such as the mathematical justification for data mixtures, the representational limits of architectures, and the optimization dynamics of alignment algorithms.
  3. 3.
    Identification of Frontier Challenges. We identify and discuss critical open questions and advanced topics that define the future of the field. By highlighting unresolved challenges, such as the theoretical limits of synthetic data self-improvement and the mathematical bounds of safety guarantees, we provide a roadmap for future scientific inquiry.
@@
The remainder of this paper is structured as follows: section 2 through section 7 provide a detailed review of the theory and mechanisms corresponding to each of the six stages. Section 8 reviews the related works, and section 9 concludes with a discussion on the future of LLM theory.
@@ 2 Data Preparation Stage
The journey of constructing an LLM begins with the data upon which it is built. The Data Preparation Stage encompasses all processes involved in collecting, cleaning, and curating the vast corpora required for training [radford2019language, albalak2023efficient, bi2025pushing]. This initial stage is arguably the most critical, as the scale, diversity, and quality of the data fundamentally define the limit of a model’s potential capabilities, including its knowledge breadth, reasoning abilities, and even its intrinsic biases [belenki-etal-2025-optimizing, muennighoff2023scaling]. While often perceived as an engineering-heavy process, the choices made during data preparation are deeply intertwined with fundamental theoretical questions about learning, generalization, and information representation. In this section, we review the theory and mechanism of the data preparation stage, from its foundational problems to the theories explaining empirical phenomena, and finally to the open questions that drive future research.
@@ 2.1 Fundamental Problems
These two questions, aiming to improve data quality and understand its impact on model performance, form the bedrock of data-centric LLM theory and mechanism. An illustration of the corresponding topics is shown in fig. 2. While classic learning theories provide the fundamental justification for scaling up datasets, their assumption of i.i.d. (independent and identically distributed) samples fails to capture the complex interplay of diversity, source mixture, and quality crucial for modern LLMs. The fundamental problem, therefore, is to extend these theories for the rich, heterogeneous, and non-i.i.d. nature of web-scale text data. To begin tackling this, we subsequently begin to review the core theories and methods in this stage.
@@ 2.2.1 Data Mixture Efficacy
A pivotal empirical finding is that performance is not merely a function of data volume, but of its heterogeneity. Models trained on a carefully curated mixture of data from multiple sources (e.g., web text, books, code, scientific articles) [liu2025rethinking] consistently outperform those trained on monolithic corpora. This observation has spurred a line of research focused on understanding and optimizing the data mixture, which has evolved along three primary axes: theoretical justification, predictive modeling, and algorithmic optimization.
Theoretical Foundations for Mixed-Data Training. The theoretical analysis is rooted in extensive classic literature on Domain Adaptation [ben2010theory, mansour2008domain, courty2016optimal]. Modern analysis for mixed data training further relies on the perspective of multi-task learning (MTL) or multi-source learning (MSL). To explain the strong generalization of highly overparameterized deep models, Zakerinia et al. [zakerinia2025low] propose a modern view based on low intrinsic dimensionality from the MTL perspective. Their key insight is that while a deep model may have a vast number of parameters, its learning process is confined to a low-dimensional manifold. Specifically, they provide a key generalization bound (Theorem 2) as follows:
@@
In eq. 1, $\mathcal{R}$ represents the true multi-task average risk , and $\hat{\mathcal{R}}$ is the empirical risk on the training data. Their core contribution is that the generalization gap (the square root term) no longer depends on the model’s vast number of original parameters, but rather on $l(E)+l_{E}(f_{1},...,f_{n})$, which represents the total compressed encoding length required to jointly encode all $n$ task models (including shared parameters $E$ and task-specific parameters $f_{i}$). This result rigorously proves theoretically that when multi-task structures are shared (allowing for shorter encoding), the model can achieve stronger generalization, even in the overparameterized state of deep learning.
Alternatively, Wang et al. [pmlr-v267-wang25eu] offer a pioneering theoretical analysis of MSL within the framework of conditional generative modeling. They establish a general distribution estimation error bound (Theorem 3.2) based on the bracketing number:
@@
In eq. 2, the average Total Variation error $\mathcal{R}_{\overline{TV}}$ is controlled by the complexity of the conditional distribution space $\mathcal{P}_{X|Y}$, as measured by its bracketing number $\mathcal{N}_{||}(\cdot)$. Their work formally addresses the question of whether it is more effective to train a single model on all sources or separate models for each one. The authors prove that when source distributions share sufficient “parametric similarity” and the model has adequate capacity, multi-source training (which results in a smaller, more constrained distribution space and thus a smaller bracketing number) is guaranteed to achieve a sharper error bound than training on sources in isolation. The theoretical advantage stems from the model’s ability to reduce the complexity of the overall distribution space it must learn.
Predictive Models for Data Mixture. Besides theoretical validation, researchers have also worked towards creating quantitative models that can predict performance based on the mix. A significant breakthrough was the introduction of “data mixing laws” [ye2024data], which establish a predictable, functional relationship between the mixing proportions of training data and the model’s validation loss on each domain. This framework allows for the a priori prediction of a model’s overall loss for any given mixture. They propose that the validation loss $L$ on a validation set composed of $K$ (potentially implicit) domains with proportions $s_{i}$, given training mixture proportions $r_{j}$ across $M$ domains, can be predicted by:
@@
This concept was further refined by BiMix [ge2024bimix], which proposes a more granular bivariate law. This model explicitly describes how two core variables—the proportion $r_{i}$ for a specific domain $i$ and the total data volume (represented by training steps ss)—jointly influence the validation loss $L_{i}$ on that specific domain:
@@
Optimization Strategies. With predictive models in place, the next logical step is to develop algorithms that automatically find the optimal mixture. Several distinct optimization strategies have emerged based on different theoretical perspectives. For instance, REGMIX [liu2025regmix] frames the problem as a regression task. It operates on the core assumption of “rank invariance,” positing that the relative superiority of a data mixture is preserved across different model scales and data volumes. Differently, UtiliMax [held2025optimizing] draws an analogy to financial portfolio optimization. It treats data sources as “assets” and seeks a mixing strategy that optimally balances three factors: the “Utility” (expected contribution) of each source, “Diversity” (to mitigate the risk of overfitting), and “Scale” (to avoid over-sampling smaller, high-quality datasets). Moreover, DoReMi [xie2023doremi] leverages min-max optimization, formulating the objective as minimizing the worst-case performance across all data domains, thereby enhancing the model’s robustness and uniformity. Another particularly powerful approach, employed by methods like DOGE [pmlr-v235-fan24e] and ScaleBIO [pan-etal-2025-scalebio], is Bilevel Optimization. This framework directly targets generalization by defining a nested objective: the outer loop optimizes the data source sampling weights to minimize validation loss, while the inner loop finds the optimal model parameters that minimize training loss given those weights.
@@ 2.2.2 Data Deduplication & Filtering
The Benefits of Data Deduplication. The investigation into deduplication’s benefits forms one line of inquiry. Lee et al. [lee-etal-2022-deduplicating] provide a foundational, empirical confirmation of its multiple advantages. This work demonstrates that deduplication directly addresses unnecessary memorization, reducing the frequency of models memorizing training text. Alternatively, Kandpal et al. [kandpal2022deduplicating] propose a core argument: data repetition is the key driver of memorization that leads to privacy risks. This work posits that privacy attacks are successful primarily because of duplicate sequences in the training data. By re-training models on sequence-level deduplicated data, they confirm that this mitigation significantly reduces privacy risks, thus establishing a causal link between data duplication and privacy vulnerabilities.
The Understanding of Deduplication. The theoretical understanding of deduplication has evolved significantly from early engineering trade-offs to more sophisticated information-theoretic concepts. Early large-scale datasets, such as The Pile [gao2020pile], recognize the importance of deduplication. However, its application was often limited by computational constraints such as only deduplicating within the noisiest subsets rather than globally. This makes deduplication an engineering-heavy compromise rather than a fully realized theoretical application. The primary bottleneck for applying deduplication theory at the trillion-scale was computational. Traditional CPU-based MinHash LSH [indyk1998approximate] implementations were too slow. This mechanism-level bottleneck was addressed by frameworks like FED [son2025fed], which introduces a reusable hash function with lower computational cost and performing end-to-end GPU parallel optimization for the entire MinHash LSH process, reducing tasks from weeks to hours. With scalability solved, the theoretical focus shifted. The RefinedWeb [penedo2023refinedweb] provided a key insight: models trained on aggressively filtered and deduplicated web data could outperform those trained on curated corpora. This suggested that data quality and information density were more critical theoretical levers than simple data source curation. The D4 [tirumala2023d4] framework further evolved this concept. It moves beyond syntactic matching (e.g., hashes) to semantic matching, leveraging pre-trained model embeddings to select a subset of documents that is both de-duplicated and semantically diverse. This mechanism demonstrated tangible performance gains, speeding up training and improving downstream accuracy. The most recent conceptual advance, SoftDedup [he-etal-2024-softdedup], addresses a theoretical flaw in “hard deduplication” methods: the risk of information loss, and further proposes a soft reweighting mechanism instead.

### v2

# Beyond the Black Box: A Survey on the Theory and Mechanism of Large Language Models

@@ 1 Introduction
The recent emergence of Large Language Models (LLMs) has marked a profound paradigm shift in the landscape of Artificial Intelligence (AI). Models such as ChatGPT (OpenAI, 2022), DeepSeek (Guo et al., 2025), Qwen (Bai et al., 2023a), Llama (Touvron et al., 2023), Gemini (Team et al., 2023), and Claude (Caruccio et al., 2024) have transcended the boundaries of traditional Natural Language Processing (NLP) (Vaswani et al., 2017a), demonstrating capabilities that impact nearly every facet of modern society. As these systems scale, they exhibit behaviors that mimic human-like reasoning (Wei et al., 2022c), sparking a global transformation in how we interact with information.
In the history of technological development, engineering triumphs are often inextricably linked to scientific innovation. However, the synchronization between theory and application is rarely instantaneous. Consider the trajectory of nuclear physics: from Einstein’s formulation of the mass-energy equivalence equation ($E=mc^{2}$) in 1905 (Einstein, 1905) to the detonation of the first atomic bomb at Los Alamos in 1945 (Rhodes, 2012), scientists and engineers traversed a forty-year journey to translate theoretical insight into physical reality (Hoddeson and Baym, 1993). A similarly extended timeline defines the current era of AI. Approximately 33 years elapsed between the proposal of the Universal Approximation Theorem (Hornik et al., 1989), which provided the mathematical assurance that neural networks could represent any continuous function, and the emergence of ChatGPT (OpenAI, 2022), the definitive proof of that potential. Looking ahead from our current vantage point, the quest for Artificial General Intelligence (AGI) necessitates a balanced synergy where continuous theoretical research and rigorous engineering implementation are recognized as equally indispensable pillars.
@@
Throughout these decades, researchers have relentlessly pursued the essence of intelligence through diverse engineering and scientific lenses. At this pivotal moment, with the empirical success of LLMs, we appear closer than ever to unveiling the nature of intelligence. Yet, a paradox persists within our current standing: despite the monumental engineering successes of LLMs, our theoretical understanding of them remains disproportionately nascent. While deep learning theory has advanced substantially (Roberts et al., 2022), the specific phenomena emerging from LLMs loom like a “dark cloud” over the field, shattering previous intuitions and challenging established statistical learning paradigms (Kaplan et al., 2020b). Consequently, we are currently forced to treat LLMs largely as “black boxes” (Linardatos et al., 2020; Zhao et al., 2024a). They function exceptionally well, yet their internal mechanisms of operation, the how and why behind their efficacy, remain elusive.
@@
The difficulty in piercing this black box stems primarily from two dimensions. First, the sheer scale of LLMs introduces unprecedented complexity (Kaplan et al., 2020b; Hoffmann et al., 2022a). With parameter counts reaching the trillions and a natural language state space that is combinatorially vast, accurately analyzing the learning dynamics and optimization landscape becomes an arduous mathematical challenge. Second, LLMs exhibit numerous “emergent” phenomena that do not appear in smaller models, such as hallucination (Xu et al., 2024b), in-context learning (ICL) (Brown et al., 2020), scaling laws (Kaplan et al., 2020b), and sudden “aha moments” during training (Guo et al., 2025). These phenomena are difficult to unify under a single theoretical framework, rendering the modeling of LLMs a fragmented endeavor. Consequently, current analyses of LLM theory and mechanisms are often scattered, isolated within specific sub-topics without a holistic view.
@@
To address this fragmentation, this survey proposes a comprehensive, lifecycle-based perspective. Following the standard LLM pipeline, we categorize the theoretical landscape into six distinct stages as illustrated in fig.˜1: the Data Preparation Stage, Model Preparation Stage, Training Stage, Alignment Stage, Inference Stage, and Evaluation Stage. By categorizing popular topics and theoretical advances into these stages, we aim to provide a structured roadmap that connects empirical observations with their underlying mechanisms.
@@
The remainder of this paper is structured as follows: section˜2 through section˜7 provide a detailed review of the theory and mechanisms corresponding to each of the six stages, and section˜9 concludes with a discussion on the future of LLM theory.
@@ 2 Data Preparation Stage
The journey of constructing an LLM begins with the data upon which it is built. The Data Preparation Stage encompasses all processes involved in collecting, cleaning, and curating the vast corpora required for training (Radford et al., 2019; Albalak et al., 2023; Bi et al., 2025). This initial stage is arguably the most critical, as the scale, diversity, and quality of the data fundamentally define the limit of a model’s potential capabilities, including its knowledge breadth, reasoning abilities, and even its intrinsic biases (Belenki et al., 2025; Muennighoff et al., 2023). While often perceived as an engineering-heavy process, the choices made during data preparation are deeply intertwined with fundamental theoretical questions about learning, generalization, and information representation. In this section, we review the theory and mechanism of the data preparation stage, from its foundational problems to the theories explaining empirical phenomena, and finally to the open questions that drive future research.
@@ 2.1 Fundamental Problems
These two questions, aiming to improve data quality and understand its impact on model performance, form the bedrock of data-centric LLM theory and mechanism. An illustration of the corresponding topics is shown in fig.˜2. While classic learning theories provide the fundamental justification for scaling up datasets, their assumption of i.i.d. (independent and identically distributed) samples fails to capture the complex interplay of diversity, source mixture, and quality crucial for modern LLMs. The fundamental problem, therefore, is to extend these theories for the rich, heterogeneous, and non-i.i.d. nature of web-scale text data. To begin tackling this, we subsequently begin to review the core theories and methods in this stage.
@@ 2.2.1 Data Mixture Efficacy
A pivotal empirical finding is that performance is not merely a function of data volume, but of its heterogeneity. Models trained on a carefully curated mixture of data from multiple sources (e.g., web text, books, code, scientific articles) (Liu et al., 2025g) consistently outperform those trained on monolithic corpora. This observation has spurred a line of research focused on understanding and optimizing the data mixture, which has evolved along three primary axes: theoretical justification, predictive modeling, and algorithmic optimization.
Theoretical Foundations for Mixed-Data Training. The theoretical analysis is rooted in extensive classic literature on Domain Adaptation (Ben-David et al., 2010; Mansour et al., 2008; Courty et al., 2016). Modern analysis for mixed data training further relies on the perspective of multi-task learning (MTL) or multi-source learning (MSL). To explain the strong generalization of highly overparameterized deep models, Zakerinia et al. (2025) propose a modern view based on low intrinsic dimensionality from the MTL perspective. Their key insight is that while a deep model may have a vast number of parameters, its learning process is confined to a low-dimensional manifold. Specifically, they provide a key generalization bound (Theorem 2) as follows:
@@
In eq.˜1, $\mathcal{R}$ represents the true multi-task average risk , and $\hat{\mathcal{R}}$ is the empirical risk on the training data. Their core contribution is that the generalization gap (the square root term) no longer depends on the model’s vast number of original parameters, but rather on $l(E)+l_{E}(f_{1},...,f_{n})$, which represents the total compressed encoding length required to jointly encode all $n$ task models (including shared parameters $E$ and task-specific parameters $f_{i}$). This result rigorously proves theoretically that when multi-task structures are shared (allowing for shorter encoding), the model can achieve stronger generalization, even in the overparameterized state of deep learning.
Alternatively, Wang et al. (2025c) offer a pioneering theoretical analysis of MSL within the framework of conditional generative modeling. They establish a general distribution estimation error bound (Theorem 3.2) based on the bracketing number:
@@
In eq.˜2, the average Total Variation error $\mathcal{R}_{\overline{TV}}$ is controlled by the complexity of the conditional distribution space $\mathcal{P}_{X|Y}$, as measured by its bracketing number $\mathcal{N}_{||}(\cdot)$. Their work formally addresses the question of whether it is more effective to train a single model on all sources or separate models for each one. The authors prove that when source distributions share sufficient “parametric similarity” and the model has adequate capacity, multi-source training (which results in a smaller, more constrained distribution space and thus a smaller bracketing number) is guaranteed to achieve a sharper error bound than training on sources in isolation. The theoretical advantage stems from the model’s ability to reduce the complexity of the overall distribution space it must learn.
Predictive Models for Data Mixture. Besides theoretical validation, researchers have also worked towards creating quantitative models that can predict performance based on the mix. A significant breakthrough was the introduction of “data mixing laws” (Ye et al., 2024a), which establish a predictable, functional relationship between the mixing proportions of training data and the model’s validation loss on each domain. This framework allows for the a priori prediction of a model’s overall loss for any given mixture. They propose that the validation loss $L$ on a validation set composed of $K$ (potentially implicit) domains with proportions $s_{i}$, given training mixture proportions $r_{j}$ across $M$ domains, can be predicted by:
@@
This concept was further refined by BiMix (Ge et al., 2024), which proposes a more granular bivariate law. This model explicitly describes how two core variables—the proportion $r_{i}$ for a specific domain $i$ and the total data volume (represented by training steps ss)—jointly influence the validation loss $L_{i}$ on that specific domain:
@@
Optimization Strategies. With predictive models in place, the next logical step is to develop algorithms that automatically find the optimal mixture. Several distinct optimization strategies have emerged based on different theoretical perspectives. For instance, REGMIX (Liu et al., 2025e) frames the problem as a regression task. It operates on the core assumption of “rank invariance,” positing that the relative superiority of a data mixture is preserved across different model scales and data volumes. Differently, UtiliMax (Held et al., 2025b) draws an analogy to financial portfolio optimization. It treats data sources as “assets” and seeks a mixing strategy that optimally balances three factors: the “Utility” (expected contribution) of each source, “Diversity” (to mitigate the risk of overfitting), and “Scale” (to avoid over-sampling smaller, high-quality datasets). Moreover, DoReMi (Xie et al., 2023) leverages min-max optimization, formulating the objective as minimizing the worst-case performance across all data domains, thereby enhancing the model’s robustness and uniformity. Another particularly powerful approach, employed by methods like DOGE (Fan et al., 2024a) and ScaleBIO (Pan et al., 2025a), is Bilevel Optimization. This framework directly targets generalization by defining a nested objective: the outer loop optimizes the data source sampling weights to minimize validation loss, while the inner loop finds the optimal model parameters that minimize training loss given those weights.
@@ 2.2.2 Data Deduplication & Filtering
The Benefits of Data Deduplication. The investigation into deduplication’s benefits forms one line of inquiry. Lee et al. (2022) provide a foundational, empirical confirmation of its multiple advantages. This work demonstrates that deduplication directly addresses unnecessary memorization, reducing the frequency of models memorizing training text. Alternatively, Kandpal et al. (2022) propose a core argument: data repetition is the key driver of memorization that leads to privacy risks. This work posits that privacy attacks are successful primarily because of duplicate sequences in the training data. By re-training models on sequence-level deduplicated data, they confirm that this mitigation significantly reduces privacy risks, thus establishing a causal link between data duplication and privacy vulnerabilities.
The Understanding of Deduplication. The theoretical understanding of deduplication has evolved significantly from early engineering trade-offs to more sophisticated information-theoretic concepts. Early large-scale datasets, such as The Pile (Gao et al., 2020), recognize the importance of deduplication. However, its application was often limited by computational constraints such as only deduplicating within the noisiest subsets rather than globally. This makes deduplication an engineering-heavy compromise rather than a fully realized theoretical application. The primary bottleneck for applying deduplication theory at the trillion-scale was computational. Traditional CPU-based MinHash LSH (Indyk and Motwani, 1998) implementations were too slow. This mechanism-level bottleneck was addressed by frameworks like FED (Son et al., 2025), which introduces a reusable hash function with lower computational cost and performing end-to-end GPU parallel optimization for the entire MinHash LSH process, reducing tasks from weeks to hours. With scalability solved, the theoretical focus shifted. The RefinedWeb (Penedo et al., 2023) provided a key insight: models trained on aggressively filtered and deduplicated web data could outperform those trained on curated corpora. This suggested that data quality and information density were more critical theoretical levers than simple data source curation. The D4 (Tirumala et al., 2023) framework further evolved this concept. It moves beyond syntactic matching (e.g., hashes) to semantic matching, leveraging pre-trained model embeddings to select a subset of documents that is both de-duplicated and semantically diverse. This mechanism demonstrated tangible performance gains, speeding up training and improving downstream accuracy. The most recent conceptual advance, SoftDedup (He et al., 2024b), addresses a theoretical flaw in “hard deduplication” methods: the risk of information loss, and further proposes a soft reweighting mechanism instead.